OpenStudy (anonymous):

Calculate the integral, if it converges. You may calculate the limit by appealing to the dominance of one function over another, or by l'Hopital's rule. dx/(4-x)^2 from 3 to 6 I have the integral as 1/(4-x) but im not sure what to do next because the hint tells me "The integral is improper at the interior of the interval; split the integral there. Determine the convergence of each part separately."

OpenStudy (anonymous):

\[\int_3^4\frac{dx}{(4-x)^2}\] and \[\int_4^6\frac{dx}{(4-x)^2}\]

OpenStudy (fibonaccichick666):

On your domain, is there any point that causes your function not to exist?

OpenStudy (anonymous):

\[\int_3^4\frac{dx}{(4-x)^2}=\lim_{l\to 4}\int_3^l\frac{dx}{(4-x)^2}\] is a start

OpenStudy (anonymous):

alright, i got -1 for that integral. but dont I also need to take it from 4 to 6? and what do i do after I get the integral for that?

OpenStudy (anonymous):

oh wait, just figured it out. it doesn't converge. haha thank you both @satellite73 and @FibonacciChick666.